# CoStruction: Conjoint radiance field optimization for urban scene reconStruction with limited image overlap

### v1

@@ 6 Additional implementation details
For both representations $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$, we use a hash grid for positional embedding [23] and spherical harmonics for directional embedding [16]. Additionally, we adopt the same appearance embedding from NeRF-W [20] to address illumination changes. Both representations utilize independent hash grids, trained separately during the training process, with identical parameters as summarized in Table 3. We also employ two additional density estimators (proposal networks), $\mathcal{F}_{\Theta}^{p_{0}}$ and $\mathcal{F}_{\Theta}^{p_{1}}$ [1], which are supervised only by the weight estimated by $\mathcal{F}_{\Theta}^{v}$, but will be used to guide the sampling for $\mathcal{F}_{\Theta}^{f}$ during initialization.
The sampling parameters used for training in our experiments are detailed in Table 4. In the refinement stage of the SDF, we replace the estimated densities provided by the proposal networks $\mathcal{F}_{\Theta}^{p_{i}},i\in\{0,1\}$ with PDF weights computed directly using $\mathcal{F}_{\Theta}^{f}$. Additionally, we slightly increase the number of samples for both coarse and fine sampling, setting them to 32 and 28, respectively, to improve the accuracy of density estimation during the refinement stage.
@@
The losses we use in CoStruction to optimize the volumetric ($\mathcal{L}_{\text{v}}$) and SDF fields ($\mathcal{L}_{\text{f}}$) are:
$\mathcal{L}_{\text{v}}=\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\hat{N}}\mathcal{L}_{\hat{N}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}}^{m},$ (11)
$\mathcal{L}_{\text{f}}=\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\hat{N}}\mathcal{L}_{\hat{N}}+\lambda_{\text{eik}}\mathcal{L}_{\text{eik}}^{m}.$ (12)
We set $\lambda_{\text{sky}}=0.01$, $\lambda_{d}=0.001$, $\lambda_{\hat{N}}=0.01$, $\lambda_{\text{sem}}=0.001$ and $\lambda_{\text{eik}}=0.1$. During the SDF refinement stage, we increase $\lambda_{\hat{N}}$ to 0.05 only on flat semantic classes and $\lambda_{d}$ to 0.1.
For Pandaset, we reduce the Eikonal loss coefficient to 0.01 during the early stage to facilitate mesh reconstruction, and then restore it to 0.1 for the final two epochs. To improve initialization on large texture-less road planes in the volumetric representation, we activate the total variation regularization on depth prediction during the first epoch for sequences 0034 and 0664 in nuScenes, and sequence 023 in Pandaset, with a weight of 1e-4.
@@
StreetSurf [12] SCILLA [9] GoF [42]-dense CoStruction (ours)
Figure 6: Additional qualitative results highlighting complex geometries. We compare our mesh to the ones generated from StreetSurf [12], SCILLA [9] and GoF [42]
@@
7 Additional details on Uncertainty Mask
One of our key contributions is to use $\mathcal{M}_{c}$ to adaptively direct the looseness of the Eikonal constraint on high frequency details and fine structures at early training stages (see our ablation study Sec. 4.2). Fig. 7 shows an example of how such a mask empirically helps divide the scene and how it evolves during training. The black areas in the mask primarily correspond to high-frequency details in the scene, such as fine structures or vegetation, while the white areas represent surface regions such as roads and buildings.
In the initial stage, $\mathcal{L}_{\text{eik}}$ is applied mainly to regions with strong ”surfaceness”, such as roads and buildings. By the end of the initialization stage, most surface regions are identified by the mask and constrained by the Eikonal loss. During the refinement stage, as the SDF learns to capture fine structures and predict consistent mesh colors, the Eikonal loss is progressively applied to these regions as well, ensuring adherence to the properties of a signed distance function.
@@ 9 Limitations
$\mathcal{M}_{c}$ $\hat{C}_{\mathcal{E}}(r)$
@@
Figure 7: An example of how $\mathcal{M}_{c}$, deduced from the mesh color prediction $\hat{C}_{\mathcal{E}}(r)$ evolves during the training stage on the KITTI-360 dataset. Where black indicates region where $\mathcal{L}_{\text{eik}}$ and $\mathcal{L}_{\text{normal}}$ are loosen during initialization.
@@
Interestingly, since 3D Gaussians are a more flexible and less constrained representation compared to SDF, GoF [42] performs better at capturing fine structures in such cases, even though it fails at large surface reconstruction under conditions of limited overlap. Applying our key concept of ”co-joint” optimization with uncertainty filtering to representations like GoF could be a promising direction for future exploration. We illustrate such cases in Fig. 8.

### v2

@@ 6 Additional implementation details
For both representations $\mathcal{F}_{\Theta}^{v}$ and $\mathcal{F}_{\Theta}^{f}$, we use a hash grid for positional embedding [23] and spherical harmonics for directional embedding [16]. Additionally, we adopt the same appearance embedding from NeRF-W [20] to address illumination changes. Both representations utilize independent hash grids, trained separately during the training process, with identical parameters as summarized in Table 4. We also employ two additional density estimators (proposal networks), $\mathcal{F}_{\Theta}^{p_{0}}$ and $\mathcal{F}_{\Theta}^{p_{1}}$ [1], which are supervised only by the weight estimated by $\mathcal{F}_{\Theta}^{v}$, but will be used to guide the sampling for $\mathcal{F}_{\Theta}^{f}$ during initialization. The sampling parameters used for training in our experiments are detailed in Table 5. In the refinement stage of the SDF, we replace the estimated densities provided by the proposal networks $\mathcal{F}_{\Theta}^{p_{i}},i\in\{0,1\}$ with PDF weights computed directly using $\mathcal{F}_{\Theta}^{f}$. Additionally, we slightly increase the number of samples for both coarse and fine sampling, setting them to 32 and 28, respectively, to improve the accuracy of density estimation during the refinement stage.
@@
The losses we use in J-NeuS to optimize the volumetric ($\mathcal{L}_{\text{v}}$) and SDF fields ($\mathcal{L}_{\text{f}}$) are:
$\mathcal{L}_{\text{f}}=\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\hat{N}}\mathcal{L}_{\hat{N}}^{u}+\lambda_{\text{eik}}\mathcal{L}_{\text{eik}}^{u}.$ (13)
We set $\lambda_{\text{sky}}=0.01$, $\lambda_{d}=0.001$, $\lambda_{\hat{N}}=0.01$, $\lambda_{\text{sem}}=0.001$ and $\lambda_{\text{eik}}=0.1$. During the SDF refinement stage, we increase $\lambda_{\hat{N}}$ to 0.05 only on flat semantic classes and $\lambda_{d}$ to 0.1. For Pandaset, we reduce the Eikonal loss coefficient to 0.01 during the early stage to facilitate mesh reconstruction, and then restore it to 0.1 for the final two epochs. To improve initialization on large texture-less road planes in the volumetric representation, we activate the total variation regularization on depth prediction during the first epoch for sequences 0034 and 0664 in nuScenes, and sequence 023 in Pandaset, with a weight of 1e-4.
@@
7 Additional details on Adaptive Thresholding
Figure 7 complements the pseudocode in Algorithm 1 by illustrating how the uncertainty threshold dynamically adjusts during training. Notably, the threshold exhibits remarkably similar convergence patterns across different datasets (KITTI-360 and Waymo). This consistency demonstrates the strong generalization capability of our self-adaptive strategy across diverse scenes.
Figure 7: The similar convergence patterns of our adaptive threshold ($\tau$) on KITTI-360 and Waymo datasets demonstrate it’s generalization capability across diverse scenes.
8 Additional details on Adaptive Relaxation
One of our key contributions is to use $\mathbb{I}_{\mu_{c}}$ to adaptively relax the Eikonal constrain and normal supervision on high frequency details and fine structures at early training stages (see Sections. 3.3). Experimentally, we set the threshold $\tau_{c}$ to 0.02 for KITTI and nuScenes, and to 0.015 for Pandaset and Waymo, as the images in these datasets are less contrastive. Fig. 11 shows an visualization of how empirically $\mathbb{I}_{\mu_{c}}$ helps to divide the scene and how it evolves during training. In the initial stage, $\mathbb{I}_{\mu_{c}}$ indicates regions with strong ”surfaceness”, such as roads and buildings. By the end of the initialization stage, most surface regions are identified by $\mathbb{I}_{\mu_{c}}$ and constrained by the Eikonal loss. During the refinement stage, as the SDF learns to capture fine structures and predict consistent mesh colors, the Eikonal loss is progressively applied to these regions as well, ensuring adherence to the properties of a signed distance function.
@@ 9 Additional experiment results
(a) StreetSurf
(b) J-NeuS (ours)
[…]
StreetSurf [12] ViiNeuS [9] GoF [43]-dense J-NeuS (ours)
[…]
@@ 10 Limitations
Interestingly, since 3D Gaussians are a more flexible and less constrained representation compared to SDF, GoF [43] performs better at capturing fine structures in such cases, even though it fails at large surface reconstruction under conditions of limited overlap. Applying our key concept of ”joint” optimization with uncertainty estimation to representations like GoF could be a promising direction for future exploration. We illustrate such cases in Figure 12.
StreetSurf [12] ViiNeuS [9] GoF [43]-dense J-NeuS (ours)
Figure 10: Additional qualitative results highlighting complex geometries. We compare our mesh to the ones generated from StreetSurf [12], ViiNeuS [9] and GoF [43]
$\mathbb{I}_{\mu_{c}}$ $\hat{C}_{\mathcal{E}}$
@@
Figure 11: An example of how the robust kernel $\mathbb{I}_{\mu_{c}}$ evolves during the training stage on the KITTI-360 dataset. Black indicates region where $\mathcal{L}_{\text{eik}}$ and $\mathcal{L}_{\hat{N}}$ are relaxed for the SDF representation.